Water is flowing into a conical vessel 18 cm deep and 10 cm across the top at the rate of 4 cm3/min. The depth of the water is "h" cm. If the rate which the surface is rising is 0.1146 cm/min, find the value of "h"?



1
Expert's answer
2022-05-27T09:33:45-0400

"D=2R=10 cm" ; "H=18cm" ; "\frac{dv}{dt}=4\frac{cm^3}{min}" ; "\frac{dh}{dt}=0.1146\frac{cm}{min}" .

Solution:

Filled volume:

"v=\frac13\pi r^2h" ; "r=h\cdot \tan\theta" ;

"v=\frac13\pi \cdot\tan^2\theta \cdot h^3" ;

"\tan{\theta}=\frac RH=\frac{D}{2H}" ;

"v=\frac13\pi \cdot(\frac{D}{2H})^2 \cdot h^3"

"\frac{dv}{dt}=\pi \cdot(\frac{D}{2H})^2 \cdot h^2\cdot\frac{dh}{dt}"

"h=\frac{2H}{D}\cdot\sqrt{\frac{1}{\pi \cdot\frac{dh}{dt}} \cdot\frac{dv}{dt}}=""\frac{2\cdot 18}{10}\cdot\sqrt{\frac{1}{\pi \cdot0.1146} \cdot4}\approx\frac{2\cdot 18}{10}\cdot\sqrt{\frac{4}{0.36}}=12\ (cm)" .

Answer: "h\approx12\ cm" .